# Accurate Unsupervised Photon Counting from Transition Edge Sensor Signals

### v2

@@ Abstract
We compare methods for signal classification applied to voltage traces from transition edge sensors (TES) which are photon-number resolving detectors fundamental for accessing quantum advantages in information processing, communication and metrology. We quantify the impact of numerical analysis on the distinction of such signals. Furthermore, we explore dimensionality reduction techniques to create interpretable and precise photon number embeddings. We demonstrate that the preservation of local data structures of some nonlinear methods is an accurate way to achieve unsupervised classification of TES traces. We do so by considering a confidence metric that quantifies the overlap of the photon number clusters inside a latent space. Furthermore, we demonstrate that for our dataset previous methods such as the signal’s area and principal component analysis can resolve up to 16 photons with confidence above $90\%$ while nonlinear techniques can resolve up to 21 with the same confidence threshold. Also, we showcase implementations of neural networks to leverage information within local structures, aiming to increase confidence in assigning photon numbers. Finally, we demonstrate the advantage of some nonlinear methods to detect and remove outlier signals.
@@ I Introduction
Photonics is a strong contender for building large-scale quantum information processing systems [1, 2, 3, 4, 5]; in many of these systems, photon number detection plays an essential role, serving as a resource for quantum advantage. These detectors can be used, for example, for the heralded generation of non-Gaussian states [6, 7, 8, 9, 10, 11, 12, 13, 14], for the sampling of classically-intractable probability distributions [15, 16, 17, 18, 19, 20] or for directly resolving multiple quanta improving the Fisher information of interferometric protocols [21, 22, 23]. The use of photon number resolving detectors provides a significant advantage as a single detector can determine the number of photons associated with a quantum state accurately [24, 25], without requiring a multiplexed network of threshold detectors with its concomitant complexity and potential inefficiency [17, 26, 27]. Transition edge sensors (TES) have been used for this task, offering resolution over a wide energy range. Resolutions up to 30 photons have been demonstrated [28], although this quantity is typically lower, on the order of 17, if more straightforward techniques are used [25].
@@
TESs exploit the superconducting phase transition of photosensitive materials to achieve an extremely sensitive calorimeter [29]. During operation, the material is cooled below its critical temperature and then current-biased to the transition region between its superconducting and normal state. In this region, the temperature increase following the absorption of a single photon leads to a measurable change in the material’s resistance [30, 31]. The resistance change is read-out using a low noise amplifier such as superconducting quantum interference devices (SQUIDs), which also enable the creation of large arrays of TES detectors via read-out multiplexing [29]. Optimized materials and coupling techniques have demonstrated efficiencies of up to 98% [32].
@@
The readout of these devices is non-trivial as the quantity one wants to determine, the energy (or the photon number for a fixed frequency), is reflected in a nonlinear fashion in the voltage signal produced by the detectors’ electronics [33]. Historically, the integral (area) of the signals has been used to assign photon numbers [25, 34]. However, distinguishing large photon numbers becomes challenging with this technique. To address this issue, linear techniques such as Principal Component Analysis (PCA) have been used [35]. A machine learning method, adapted from the K-means algorithm to account for the Poissonian statistics of laser sources, has also been developed [36]. However, these methods’ simplicity or assumptions can limit their performance or usability for model-free photon number detection and when measuring non-classical sources, which typically do not have Poisson photon-number statistics.
@@
With the increased popularity of machine learning in the field of signal processing [37] and quantum systems [38], one might naturally ask whether employing more sophisticated methods could lead to enhanced resolution of photon numbers. In this work, we answer this question by assessing the performance of multiple techniques for photon number classification using TES signals. We do so by considering a confidence metric that quantifies the overlap of the photon number clusters inside a latent space. We demonstrate that for our dataset previous methods such as the signal’s area and PCA can resolve up to 16 photons with confidence above $90\%$ while nonlinear techniques can resolve up to 21 with the same confidence threshold. Furthermore, we also showcase implementations of neural networks to leverage information within local structures, aiming to increase confidence in assigning photon numbers. Finally, we demonstrate the advantage of some nonlinear methods to detect and remove outlier signals.
@@
Our manuscript is structured as follows: in the next section, Sec. II, we formulate the problem of photon-number discrimination in the general setting of unsupervised classification and dimensionality reduction. Next, in Sec. III, we offer a brief overview of the methods used to compute similarities between signals and how we distinguish signals that belong to the different photon number classes. We present our results in Sec. IV using experimental data, followed by a discussion of the use cases of the described methods in Sec. V
@@ II.2 Dimensionality Reduction
[…]
@@
We illustrate the process in Fig. 1 by transforming the TES signals (Fig. 1a) into one-dimensional samples presented in Fig. 1b. This low dimensional space is visualized using a kernel density estimation of the latent space (Gaussian kernel) [39]. From the position of the samples in the latent space (never considering the density estimation in the computation) it is possible to find regions most likely to describe a photon number $n\in\{0,1,...,8\}$, we discuss this step in Sec. III.4. Finally, from this interpretation of the low-dimensional space, a photon number can be assigned to every sample (Fig. 1c). The regions of high density in Fig. 1b are called clusters and are associated with photon numbers. We note that clusters can be defined using other heuristics like neighbour distances.
@@ III Methods
With experimental motivations, we consider the properties and use cases of dimensionality reduction techniques. To do so, the methods are divided into three categories based on their characteristics: basic feature, non-predictive, and predictive.
@@ III.1 Basic features
The methods in this category rely on some feature with physical significance, and their latent space represents the value of this feature. These methods are fast to compute due to their simplicity and can be combined with noise filtering to increase resolution [25].
@@ III.1.1 Maximum Value
The maximum value of the signals has been used in some cases for photon number resolution [25]. For experiments that only require the measurement of low photon numbers, sufficient information is found in the maximum value. For high enough photon numbers, the traces reach a plateau and the maximum value no longer gives information [33].
@@ III.1.2 Area
TES pulse area relates non-linearly to the energy absorbed by the sensor and therefore can be used for dimensionality reduction [25]. The area is sensitive to noise outside the pulse, hence filtering and background rejection are used in some cases to increase the performance of this method. To offer a fair representation of this technique, a Butterworth filter is applied to the signals and a threshold is introduced to reduce the influence of noise. Following existing work, the threshold is defined above the noise distribution in the flat region of the TES signals (where only vacuum is detected) [25].
@@ III.2 Non-predictive methods
The method t-SNE is non-predictive and attempts to create a low-dimensional representation of the data by organizing all the samples in a low-dimensional space. The position of the samples is assigned using a gradient descent by minimizing the Kullback-Leibler divergence ($\mathrm{KL}$)
@@
In the $\mathrm{KL}$ divergence, $p_{ij}$ represents joint probabilities that describe the similarities between high-dimensional samples $x_{i}$ and $x_{j}$ and is the $q_{ij}$ joint probabilities for low-dimensional samples $y_{i}$ and $y_{j}$ [41]. The high-dimensional joint probabilities are set to be symmetric conditional probabilities defined as
@@
$p_{j|i}=\frac{\exp\left[-\tfrac{1}{2}||x_{i}-x_{j}||^{2}/\sigma_{i}^{2}\right]}{\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{u}\exp\left[-\tfrac{1}{2}||x_{i}-x_{k}||^{2}/\sigma_{i}^{2}\right]},$ (3)
where $||x||=(\sum_{i}x_{i}^{2})^{1/2}$ represents the Euclidean norm. In low-dimensional space, the joint probabilities are given by Student t-distribution
$q_{ij}=\frac{(1+||y_{i}-y_{j}||^{2})^{-1}}{\sum_{k=1}^{u}\sum_{\begin{subarray}{c}l=1\\ l\neq k\end{subarray}}^{u}(1+||y_{k}-y_{l}||^{2})^{-1}}.$ (4)
@@
This parameter, initially introduced in speech recognition, is user-defined and is often described as an effective number of neighbours [42]. The intuition behind this value is that the variance of each Gaussian in the high dimensional space is tuned to have a tail with a limited number of relevant neighbours. This means neighbours outside the effective range of the Gaussian will have similarity values considerably smaller.
@@ III.3.3 Non-Negative Matrix Factorization (NMF)
$L(\bm{X},\bm{Y},\bm{H})=||\bm{X}-\bm{Y}\bm{H}||_{\text{Frob}}^{2}.$ (17)
The Frobenius norm is a matrix norm defined for a matrix $\bm{A}$ with elements $a_{ij}$ as $||\bm{A}||_{\text{Frob}}=(\sum_{ij}|a_{ij}|^{2})^{1/2}$. It can be shown that the optimization of the Frobenius norm is equivalent to the maximum likelihood estimate of $\bm{X}$ without Gaussian noise [50]. Additionally, we test NMF optimization using the $\mathrm{KL}$ divergence, where the loss function becomes
@@ III.3.4 Neural Networks
Neural networks have the potential to reproduce a wide variety of operations in a numerical structure that can be used efficiently to process large amounts of data. To quickly apply UMAP and t-SNE on new data we use parametric implementations of these methods using neural networks. The main principle behind these parametric implementations is to constrain the embedding to transformations done through a neural network. In other words, a neural network is trained to optimize the $\mathrm{KL}$ divergence in the case of t-SNE and the cross-entropy in UMAP. By applying this constraint during training, we create a neural network that considers local structures and behaves similarly to t-SNE and UMAP. At this stage, new data can be embedded at an efficiency restricted by the complexity of the neural network architecture.
More details about the neural network architecture and the training process are provided in Appendix A.
@@ III.5 Number of clusters
The criteria assign a score given a number of clusters $K$, a likelihood function $\mathcal{L}(\theta)$, and a total number of data points $u$. By sweeping the number of clusters used in some models, these criteria give a way to find a balance between the number of clusters and the likelihood. In our case, the likelihood of the Gaussian mixture model is used to evaluate the information scores. The general idea of these criteria is to negatively score the number of clusters, considering it is always possible to overfit the data with more clusters. In other words, a model with more clusters can always achieve a higher or equal likelihood than a model with fewer clusters. The point of diminishing return is given by the “elbow” of the $\mathrm{AIC}$ and $\mathrm{BIC}$ when evaluating the criteria as a function of the number of clusters. After this point, the additional clusters mostly overfit the data.
The Silhouette score is also used with the information criteria to evaluate the number of clusters [51]. Since similar results are found with this method, the details are not described here.
@@ III.6 Quality Assessment
Assessing the performance of dimensionality reduction techniques in an unsupervised setting is difficult since the ground truth is unknown. To tackle this task, we quantify cluster separation. To improve the performance evaluation it is also important to understand that the problem is not completely unsupervised considering photon sources used to generate samples follow known distributions. We include this knowledge of photon number distributions as an additional validation to cluster separation evaluation in the confidence metric (Sec. III.6.1).
@@ III.6.1 Confidence
We consider the probability density of photon events can be approximated from the sample’s distribution in the latent space following the Gaussian mixture model. Following previous work [35], the confidence $C_{n}$ is used as a performance metric for the resolution of photon numbers in a latent space following,
@@ III.7 Datasets
Table 1: Number of samples $u$, number of time steps $t$ and photon number distribution for both the training and testing portion of all the datasets used in this work. For cases where the photon number distribution is engineered to resemble a goal distribution, the blue bars represent the expected photon number distribution for a mixture of Poisson distribution and the yellow bars are the goal distributions used to fit the weights $w_{\langle n\rangle}$.
Name Number ($u$) Size ($t$) Distribution Reference
Synthetic Uniform
Train : 30 5503055030\,55030 550
Test : 30 5503055030\,55030 550
$350$ [Uncaptioned image] [33, 52]
Synthetic Geometric
Train : 57 0205702057\,02057 020
Test : 57 0205702057\,02057 020
$350$ [Uncaptioned image] [33, 52]
Synthetic Large
Train : 550 000550000550\,000550 000
Test : 550 000550000550\,000550 000
$200$ [Uncaptioned image] [53]
Noise Test : 200 000200000200\,000200 000 $50$ [Uncaptioned image] [53]
[…]
Instead of directly using these distributions, we construct two synthetic datasets (made of real traces) that follow a close-to-uniform and close-to-geometric distributions $P(n)$. These datasets are labelled as Synthetic Uniform and Synthetic Geometric in Table 1. Furthermore, for both of these datasets, a training and testing set were generated. Considering randomly selecting a portion of the samples in each experiment is equivalent to varying the weight $w_{{\langle n\rangle}}$ of a given Poisson distribution $P_{{\langle n\rangle}}(n)$ inside a mixture of Poisson distributions. The total expected distribution $P(n)$ can be described by
@@
and where $\bar{N}$ is the set of available mean photon numbers $\langle n\rangle$. With this construction, the expected photon number distribution is a mixture of Poisson distributions shown in Table. 1. The choice of a uniform distribution is motivated by the desire to make the labelling task difficult by maximizing the distribution’s entropy. In other words, for every sample in a perfectly uniform distribution, the method would have equal chances of guessing every class. The choice of testing a geometric distribution comes from the desire to precisely measure thermal optical sources that follow a geometric photon number distribution. Also, distributions with a long tail can be difficult to process for certain methods since fewer examples are present in some classes (imbalanced dataset).
[…]
Finally, we also make use of a dataset labelled Noise, in Sec. V.5, for this dataset, $u=200000$ TES signals were produced by detecting light generated by an integrated optical parametric oscillator (OPO) pumped below threshold using a pulsed-carved continuous wave laser, as in Ref. [54]. The OPO generated signal photons following a quasi-thermal distribution. In addition, noise photons from the pump leaked into the detected mode due to imperfect pulse carving and filtering. These noise photons were generated at random times relative to the signal photons. All datasets are summarized in Table 1.

### v3

# Accurate Unsupervised Photon Counting from Transition-Edge Sensor Signals

@@ Abstract
We compare methods for signal classification applied to voltage traces from transition-edge sensors (TES) which are photon-number resolving detectors fundamental for accessing quantum advantages in information processing, communication and metrology. We quantify the impact of numerical analysis on the distinction of such signals. Furthermore, we explore dimensionality reduction techniques to create interpretable and precise photon-number embeddings. We demonstrate that the preservation of local data structures of some nonlinear methods is an accurate way to achieve unsupervised classification of TES traces. We do so by considering a confidence metric that quantifies the overlap of the photon-number clusters inside a latent space. Furthermore, we demonstrate that for our dataset previous methods such as the signal’s area and principal component analysis can resolve up to 16 photons with confidence above $90\%$ while nonlinear techniques can resolve up to 21 with the same confidence threshold. Also, we showcase implementations of neural networks to leverage information within local structures, aiming to increase confidence in assigning photon numbers. Finally, we demonstrate the advantage of some nonlinear methods to detect and remove outlier signals.
@@ I Introduction
Photonics is a strong contender for building large-scale quantum information processing systems [1, 2, 3, 4, 5]; in many of these systems, photon-number detection plays an essential role, serving as a resource for quantum advantage. Photonic architectures often encode information in continuous variables or in multi-photon states, where precise knowledge of photon number is critical for state preparation, measurement, and error correction. For example, photon-number-resolving detectors have been used for the generation of non-Gaussian states [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], for the sampling of classically-intractable probability distributions [17, 18, 19, 20, 21, 22], or for directly resolving multiple quanta, thereby improving the Fisher information in interferometric protocols [23, 24, 25]. In such protocols, quantum states of light such as N00N or Holland-Burnett states can, in principle, achieve a precision scaling of $\sqrt{N}$ (for $N$ detected photons), surpassing classical limits. However, this quantum advantage is highly sensitive to loss, making efficient photon-number resolution particularly valuable. The use of photon-number resolving detectors offers a significant advantage, as a single device can accurately determine the number of photons in a quantum state [26, 27], thereby eliminating the need for a demultiplexed network of threshold detectors, which adds complexity and can reduce overall efficiency [19, 28, 29]. Transition-edge sensors (TES) have been used for this task, offering resolution over a wide energy range. Resolutions up to 30 photons have been demonstrated [30], although this quantity is typically lower, on the order of 17 [27].
@@
TESs exploit the superconducting phase transition of photosensitive materials (illustrated in Fig. 1(a)) to achieve an extremely sensitive calorimeter [31]. During operation, the material is cooled below its critical temperature and then current-biased to the transition region between its superconducting and normal state. In this region, the temperature increase following the absorption of a single photon leads to a measurable change in the material’s resistance [32, 33]. The resistance change is read-out using a low noise amplifier such as superconducting quantum interference devices (SQUIDs), which also enable the creation of large arrays of TES detectors via read-out multiplexing [31]. Optimized materials and coupling techniques have demonstrated system efficiencies of up to 98% [34].
@@
The readout of these devices is non-trivial as the quantity one wants to determine, the energy (or the photon number for monochromatic light), is reflected in a nonlinear fashion in the voltage signal produced by the detectors’ electronics [35]. Historically, the integral (area) of the signals has been used to assign photon numbers [27, 36]. However, distinguishing large photon numbers becomes challenging with this technique. To address this issue, linear techniques such as Principal Component Analysis (PCA) have been used [37]. A machine learning method, adapted from the K-means algorithm to account for the Poissonian statistics of laser sources, has also been developed [38]. However, these methods’ simplicity or assumptions can limit their performance or usability for model-free photon-number detection and when measuring non-classical sources, which typically do not have Poisson photon-number statistics.
@@
With the increased popularity of machine learning in the field of signal processing [39] and quantum systems [40], one might naturally ask whether employing more sophisticated methods could lead to enhanced resolution of photon numbers. In this work, we answer this question by assessing the performance of multiple techniques for photon number classification using TES signals. We do so by considering a confidence metric that quantifies the overlap of the photon-number clusters inside a latent space. We demonstrate that for our dataset previous methods such as the signal’s area and PCA can resolve up to 16 photons with confidence above $90\%$ while nonlinear techniques can resolve up to 21 with the same confidence threshold. Furthermore, we also showcase implementations of neural networks to leverage information within local structures, aiming to increase confidence in assigning photon numbers. Finally, we demonstrate the advantage of some nonlinear methods to detect and remove outlier signals.
@@
Our manuscript is structured as follows: in the next section, Sec. II, we formulate the problem of photon-number discrimination in the general setting of unsupervised classification and dimensionality reduction. Next, in Sec. III, we offer a brief overview of the methods used to compute similarities between signals and how we distinguish signals that belong to the different photon-number classes. We present our results in Sec. IV using experimental data, followed by a discussion of the use cases of the described methods in Sec. V
@@ II.2 Dimensionality Reduction
(a) Circuit diagram of a typical transition-edge sensor detection scheme. The circuit can change slightly from one implementation to the other, the illustrated circuit is based on Ref. [27]. The superconducting phase transition of the TES is illustrated through the sharp variation of resistance as a function of temperature, giving the TES extremly sensitive energy resolution [41].
(b) Example of multiple voltage outputs measured by an oscilloscope to construct a dataset $\bm{X}$ with u=1 024u=1\,024 raw TES traces of size $t=100$.
(c) The dataset $\bm{X}$ is transformed into $\bm{Y}$ which has a single dimension ($r=1$), here plotted using a kernel density estimation [42]. The dimensionality reduction technique (maximum value of the signals in this case) creates a low-dimensional space where signal features become apparent. Each peak is a cluster that represents the underlying dominant feature of the signals: the photon numbers.
(d) Clusters in the latent space are assigned a photon number $n\in\{0,1,...,8\}$. This is done by dividing the space in regions most likely associated to specific photon numbers (see Sec. III.6.1). From labelled samples, a photon-number distribution can be generated.
Figure 1: Steps associated with the photon number detection process going from the device in (1(a)) and the expected output signals in (1(b)) to the abstract space describing similarities between samples in (1(c)) and the final photon number distribution associated with an experiment in (1(d)).
@@
We illustrate the process in Fig. 1 by transforming the TES signals (Fig. 1(b)) into one-dimensional samples presented in Fig. 1(c). This low dimensional space is visualized using a kernel density estimation of the latent space (Gaussian kernel) [42]. From the position of the samples in the latent space (never considering the density estimation in the computation) it is possible to find regions most likely to describe a photon number $n\in\{0,1,...,8\}$. We discuss this step in Sec. III.4. Finally, from this interpretation of the low-dimensional space, a photon number can be assigned to every sample (Fig. 1(d)). The regions of high density in Fig. 1(c) are called clusters and are associated with photon numbers. We note that clusters can be defined using other heuristics like neighbour distances.
@@ III Methods
With experimental motivations, we consider the properties and use cases of dimensionality reduction techniques. To do so, the methods are divided into three categories based on their characteristics: basic features, non-predictive, and predictive.
@@ III.1 Basic Features
The methods in this category rely on some feature with physical significance, and their latent space represents the value of this feature. These methods are fast to compute due to their simplicity and can be combined with noise filtering to increase resolution [27]. To offer a fair representation of these techniques, we apply a Butterworth low-pass filter to the signals. We do not explore alternative representation of the traces by strongly reducing the frequency content of the signals. The filtering is only used in this case to remove high frequency electrical noise and preserve the original shape of the signals.
@@ III.1.1 Maximum Value
The maximum value of the signals has been used in some cases for photon-number resolution [27]. For experiments that only require the measurement of low photon numbers, sufficient information is found in the maximum value. For high enough photon numbers, the traces reach a plateau and the maximum value no longer gives information [35].
@@ III.1.2 Area
TES pulse area relates non-linearly to the energy absorbed by the sensor and therefore can be used for dimensionality reduction [27]. The area is sensitive to noise outside the pulse, hence filtering and background rejection are used in some cases to increase the performance of this method. In this work, we only filter the traces and set a threshold to reduce the influence of noise. Following existing work, the threshold is defined above the noise distribution in the flat region of the TES signals (see time steps 80 to 100 in Fig. 1(b), where only vacuum is detected) [27].
@@ III.2 Non-Predictive Methods
The method t-SNE is non-predictive and attempts to create a low-dimensional representation of the data by organizing all the samples in a smaller space. The position of the samples is assigned using a gradient descent by minimizing the Kullback-Leibler divergence ($\mathrm{KL}$)
@@
In the $\mathrm{KL}$ divergence, $p_{ij}$ represents the joint probabilities that describe the similarities between high-dimensional samples $x_{i}$ and $x_{j}$, and $q_{ij}$ are the joint probabilities for low-dimensional samples $y_{i}$ and $y_{j}$ [44]. The high-dimensional joint probabilities are set to be symmetric conditional probabilities defined as
@@
[…]
where $x||=(\sum_{i}x_{i}^{2})^{1/2}$ represents the Euclidean norm. In low-dimensional space, the joint probabilities are given by the Student t-distribution
[…]
@@
Perplexity, initially introduced in the field of speech recognition, is user-defined and is often described as an effective number of neighbours [45]. The intuition behind this value is that the variance of each Gaussian in the high dimensional space is tuned to have a tail with a limited number of relevant neighbours. This means neighbours outside the effective range of the Gaussian will have similarity values considerably smaller.
@@ III.2.3 Isometric Mapping (Isomap)
Since this technique does not offer significant advantages (see Sec. IV) it is not discussed in further details.
@@ III.3.3 Non-Negative Matrix Factorization (NMF)
[…]
The Frobenius norm is a matrix norm defined for a matrix $\bm{A}$ with elements $a_{ij}$ as $\bm{A}||_{\text{Frob}}=(\sum_{ij}|a_{ij}|^{2})^{1/2}$. It can be shown that the optimization of the Frobenius norm is equivalent to the maximum likelihood estimate of $\bm{X}$ without Gaussian noise [53]. Additionally, we test NMF optimization using the $\mathrm{KL}$ divergence, where the loss function becomes
@@ III.3.4 Neural Networks
Neural networks have the potential to reproduce a wide variety of operations in a numerical structure that can be used efficiently to process large amounts of data. To quickly apply UMAP and t-SNE on new data, we use parametric implementations of these methods using neural networks. The main principle behind these parametric implementations is to constrain the embedding to transformations done through a neural network. In other words, a neural network is trained to optimize the $\mathrm{KL}$ divergence in the case of t-SNE and the cross-entropy in UMAP. By applying this constraint during training, we create a neural network that considers local structures and behaves similarly to t-SNE and UMAP. At this stage, new data can be embedded at an efficiency restricted by the complexity of the neural network architecture.
For both parametric implementations of t-SNE and UMAP, we use a simple feed-forward neural network defined as a series of blocks containing linear layers with ReLU activation functions followed by a batch normalization step. The results presented in this work use a neural network with 4 blocks, where each linear layer contains 300 inputs and outputs. While more complex architectures could be used for this task, we find that even an elementary neural network can achieve this task accurately, resulting in fast data transformation adequate for real-time processing.
We note that we use the same neural network to predict data for both close-to-uniform and close-to-geometric cases. This is done to train the neural network on a balanced dataset, and in this process we guarantee that the neural network is never trained on test data. Using different distributions for the training step is not advantageous to parametric methods, since the close-to-uniform dataset contains fewer samples than in the close-to-geometric case.
@@ III.4 Clustering
After computing the Gaussian mixture, photon numbers are assigned based on the average area of the samples inside each cluster. The mean area provides a reference for the cluster ordering, since a signal’s area grows with energy and, consequently, the number of photons. This task is unambiguous because the average areas are well separated, unlike the individual samples.
@@ III.5 Number of Clusters
The criteria assign a score given a number of clusters $K$, a likelihood function $\mathcal{L}(\theta)$, and a total number of data points $u$. By sweeping the number of clusters used in some models, these criteria give a way to find a balance between the number of clusters and the likelihood. In our case, the likelihood of the Gaussian mixture model is used to evaluate the information scores. The general idea of these criteria is to negatively score the number of clusters, considering it is always possible to overfit the data with more clusters. In other words, a model with more clusters can always achieve a higher or equal likelihood than a model with fewer clusters. The point of diminishing return is given by the “elbow” of the $\mathrm{AIC}$ and $\mathrm{BIC}$ when evaluating the criteria as a function of the number of clusters, see Fig. 2. After this point, the additional clusters mostly overfit the data.
Figure 2: Akaike and Bayesian information criterion (AIC and BIC) as a function of the number of clusters in the context of a maximum value latent space on the Synthetic Uniform dataset. The point of diminishing returns gives an estimate for the number of clusters in the latent space. The Silhouette score is also plotted, where the maximum value gives an estimate for the number of clusters.
The maximum Silhouette score (see Fig. 2) is also used along the information criteria to evaluate the number of clusters [55]. Since similar results are found with this method, the details are not described here.
@@ III.6 Quality Assessment
Assessing the performance of dimensionality reduction techniques in an unsupervised setting is difficult since the ground truth is unknown. To tackle this task, we quantify cluster separation. To improve the performance evaluation it is also important to understand that the problem is not completely unsupervised considering photon sources used to generate samples follow known distributions. We include this knowledge of photon-number distributions as an additional validation to cluster separation evaluation in the confidence metric (Sec. III.6.1).
@@ III.6.1 Confidence
We consider the probability density of photon events can be approximated from the sample’s distribution in the latent space following the Gaussian mixture model. Following previous work [37], the confidence $C_{n}$ is used as a performance metric for the resolution of photon numbers in a latent space, following,
@@ III.7 Datasets
(a) Synthetic Uniform : Composed of u=30 550u=30\,550 train and test samples of size $t=350$, generated using an attenuated 1550 nm coherent source [35, 56].
(b) Synthetic Geometric : Composed of u=57 020u=57\,020 train and test samples of size $t=350$, generated using an attenuated 1550 nm coherent source [35, 56].
(c) Synthetic Large : Composed of u=550 000u=550\,000 train and test samples of size $t=100$, generated using a 1530nm coherent source [57].
(d) Noise : Composed of u=1 000 000u=1\,000\,000 samples in both channels of size $t=50$, generated using twin beams at 1520 nm and 1580 nm. Noise at different wavelengths is also present in the data and is filtered out using UMAP in the plotted distribution (not characterized) [57].
Figure 3: Compilation of all datasets used in this research. For cases where the photon-number distribution is engineered to resemble a goal distribution, the blue bars represent the expected photon-number distribution for a mixture of Poisson distribution and the yellow bars are the goal distributions used to fit the weights $w_{\langle n\rangle}$.
Experimental data from previous work at the National Institute of Standards and Technology (NIST) is used to benchmark the different techniques in this work [35]. The original dataset was generated by progressively attenuating a coherent source from 29dB to 7dB, leading to 24 datasets each containing u=20 480u=20\,480 signals and t=8 192t=8\,192 time steps. This results in datasets that each have Poisson photon-number distributions and mean photon number $\braket{n_{1}}=2.26$ to $\braket{n_{24}}=7.08\times 10^{6}$. These values were independently measured using a calibrated photodetector.
Instead of directly using these distributions, we construct two synthetic datasets (made of real traces) that follow a close-to-uniform and close-to-geometric distribution $P(n)$. These datasets are labelled as Synthetic Uniform and Synthetic Geometric in Fig. 3. Furthermore, for all datasets, a training and testing set were generated by taking respectively the first and second half of the files for every attenuation. This simulates a calibration and measurement stage but effectively does not change the results. Considering randomly selecting a portion of the samples in each experiment is equivalent to varying the weight $w_{{\langle n\rangle}}$ of a given Poisson distribution $P_{{\langle n\rangle}}(n)$ inside a mixture of Poisson distributions. The total expected distribution $P(n)$ can be described by
@@
and where $\bar{N}$ is the set of available mean photon numbers $\langle n\rangle$. With this construction, the expected photon-number distribution is a mixture of Poisson distributions shown in Fig. 3. The choice of a uniform distribution is motivated by the desire to make the labelling task difficult by maximizing the distribution’s entropy. In other words, for every sample in a perfectly uniform distribution, the method would have equal chances of guessing every class. The choice of testing a geometric distribution comes from the desire to precisely measure thermal optical sources that follow a geometric photon-number distribution. Also, distributions with a long tail can be difficult to process for certain methods since fewer examples are present in some classes (imbalanced dataset).
Additionally, these expected distributions are used as $P(n)$ in the computation of the confidence. The predictive methods are trained with the training set, and the analysis of performance metrics is done by feeding the test set to the trained methods. In the case of non-predictive and basic feature methods, the test set is directly used. The training and test datasets contain a total of u=30 550u=30\,550 traces of size $t=350$ (first $350$ values of the $8192$ available time steps). We note that most of the weights $w_{\braket{n}}$ are set to zero because of the number of available Poisson distributions in the desired photon number range is small, making the synthetic distribution not perfectly uniform (see top row in Fig. 3).
To validate a hypothesis discussed in Sec. V.2 we also use a larger dataset named Synthetic Large that was created using signals generated by TESs at the National Research Council Canada (NRC) in Ottawa [57]. The data was generated by tuning the attenuation of a 1530 nm laser and measuring u=100 000u=100\,000 signals for each of these coherent sources.
Finally, we also make use of a dataset labelled Noise, in Sec. V.5, the light is generated by an integrated optical parametric oscillator (OPO) pumped below threshold using a pulsed-carved continuous wave laser, as in Ref. [58]. The OPO generated correlated 1520 nm and 1580 nm photons following twin beams photon statistics. In addition, noise photons from the pump leaked into both modes due to imperfect pulse carving and filtering. These photons were generated at random times relative to the signal photons, which reduced the photon correlation strength between the two modes. Tracing over one of these modes, we expect quasi-thermal photon statistics. For this dataset, u=1 000 000u=1\,000\,000 TES signals were recorded from each detector channel. All datasets are summarized in Fig. 3.